# Systematic Generalization in Language Models Scales with Information Entropy

### v1

@@ 1 Introduction
Human language is characterized by its combinatorial properties in syntax and semantics (Hadley, 1994). For example, when somebody knows what it means for Person A to see Person B, they can readily generalize to what it means for Person B to see Person A, without having seen examples of language use where Person B is used in the subject position (Fodor and Pylyshyn, 1988). This type of bidirectional systematic generalization has long been considered a key feature of human language understanding (Chomsky, 1957).
@@
Early critics of connectionism, such as Fodor and Pylyshyn (1988), argued that systematic generalization could not be implemented using continuous representations. Today, over thirty years after this critique was initially presented, researchers agree that achieving robust compositional behavior in neural networks remains an open problem, as shown by a recent survey by McCurdy et al. (2024). Indeed, research continues to show that current models struggle with systematic generalization (Lake and Baroni, 2018; Hupkes et al., 2020; Dziri et al., 2024; Wold et al., 2024; Mirzadeh et al., 2024). Although existing works have overcome some of these limitations by implementing strong priors (Gordon et al., 2019; Lake and Baroni, 2023), the reason for this apparent lack of generalization remains unclear: are limitations caused by inherent architectural constraints, or due to distributional properties of the data used to train models?
@@
Figure 2: A schematic overview of the two approaches to increasing the entropy level $\mathrm{H}$. Top: Vertical scaling by distribution mixing. Bottom: Horizontal scaling is done by incrementing the support of the distribution.
In this paper, we ask the following question: Can neural networks display systematic out-of-distribution generalization without any built-in priors? To answer this, we take a data-centric approach. Specifically, we show how common architectures’ degree of systematic generalization relates to the information entropy of the distribution of component parts within training samples.
Contrary to previous work, our findings show that popular sequence-to-sequence models used in NLP are capable of systematic generalization—even without built-in architectural priors. However, this is only the case under high entropy. Through a series of sequence-to-sequence tasks based on a modified version of the SCAN (Lake and Baroni, 2018) grammar, we show that model performance scales with the entropy of the training dataset, as illustrated in Figure 1, thereby creating a link between systematic generalization and information efficiency.
@@ 2 Background
Recent work by McCurdy et al. (2024, p. 9324) show that there is high agreement among researchers on the following definition of compositional behavior:
@@
“When a model receives an input I that humans conceive as composed of component parts, if the model produces correct outputs for those parts (in isolation or in other combinations), then it will also produce a correct output for I.”
@@ 2.1 Systematic Generalization
Following earlier works, we will refer to the systematic generalization abilities of a model as the systematicity of that model. Our view of systematicity is based on Hadley (1994), who in contrast to earlier work by Fodor and Pylyshyn (1988) distinguished between degrees of systematicity. In the following paragraphs, we define these degrees under the assumption that we are considering a system tasked to generalize from a training corpus to a held-out test corpus.
Weak systematicity.
@@
Quasi-systematicity.
A system displays quasi-systematicity if it (i) displays weak systematicity and (ii) can process novel sentences containing embedded sentences, [^1] In this context, an embedded sentence can be thought of as a syntactic constituent in the overall sequence. such that both the overall sentence and the embedded sentences are syntactically isomorphic to a sentence in the training data, and (iii) for every successfully processed sentence, the component parts in the embedded sentence occur in the same syntactic position in a non-embedded sentence in the training-corpora. This means that a system displays quasi-systematicity if it needs to see a component part at a specific syntactic position in a non-embedded sentence if it is to generalize to that same component when used in an embedded sentence.
Strong systematicity.
Finally, a system displays strong systematicity if it (i) displays weak systematicity and (ii) it can process novel embedded sentences containing previously learned component parts that occur in syntactic positions where they do not appear in the training data, neither in embedded sentences nor in non-embedded sentences, and (iii) a significant fraction of the possible combinations of component parts and their permissible syntactic position should be left out, based on the observation that humans learn grammars based on largely incomplete data Pinker (1989).
@@
In this work, we will target a relaxed version of strong systematicity. We relax requirement (ii) from the original definition so that we only work with samples that have embedded sentences. As we show in the following section, this makes it easier to quantify the level of systematicity required for a given generalization problem, which is the main focus of this work. Furthermore, we also relax (ii) so that it allows component parts in the test set to occur at the same syntactic position in some embedded sentence in the train set, but not in an embedded sentence that occur at the same position in the overall sequence. In order to compare models with and without compositional priors from the literature, we base our experiments on the SCAN grammar (Lake and Baroni, 2018). This grammar has few permissible syntactic positions for each part of speech, making the original requirement from Hadley (1994) too restrictive with respect to the overall sample size needed by neural networks. In the following section, we formalize our framework and elaborate on these choices.
@@ 3 Measuring Systematicity
In this work, we will demonstrate that neural networks are able to generalize to embedded sentences where component parts occur at novel syntactic positions in the overall sequence. In our experiments, we will use a modified version of the SCAN grammar. We construct datasets where each dataset sample contains an input sequence $x\in X$ and an output sequence $y\in Y$. We discuss the specifics of this dataset generation procedure, including details about the context-free grammar (CFG) that generates $x$, in Section 3.2. In this section, we define a framework for measuring the degree of systematicity displayed by a systematic generalization problem.
Throughout this paper, $x$ will comprise two embedded sentences conjoined by a conjunctive phrase. [^2] In contrast to the original SCAN, we do not have non-embedded sentences (e.g. sequences that do not have a conjunctive phrase). Consequently, $x$ can be modeled as a triple, with a conjunction $\operatorname{c}\in C$, and the two embedded sentences $e_{1},e_{2}\in\Sigma^{*}$ as arguments, $x=(e_{1},\operatorname{c},e_{2})$. The component parts of both embedded sentences are drawn from the same vocabulary $\Sigma=V\cup A$, where $V$ is a set of verbs and $A$ is a set of adverbial phrases, e.g: jump, run $\in V$, twice $\in A$, and and $\in C$. There exists a deterministic procedure $\varphi$ for transforming $x$ into $y$:
@@
The goal of the described task is to learn a function $f$ that approximates $\varphi$, which requires modeling the conditional distribution of output sequences given input sequences in the training data, i.e. $p^{\text{train}}(y\mid x)$. See Table 1 for more examples generated by our modified SCAN grammar. Since we are interested in systematic generalization, we need to also define the distribution of component parts in $e_{1}$ and $e_{2}$, and we are specifically interested in the case where the distribution of $e_{1}$ and $e_{2}$ is different in the training and test data, but they share a similar distribution when marginalized over position, which in this case refers to which side of the conjunctive phrase they appear in.
In SCAN (Appendix A), $e_{1}$ and $e_{2}$ always contain one verb in $V$ and zero or more adverbial phrases from $A$ that control how many times that verb should be repeated, and in what direction. To create a systematic generalization problem that adheres to the definition given in Section 2.1, it is necessary to set the support of $e_{1}$ and $e_{2}$ to be different in the training and test distributions. For simplicity, we limit our study to focus on the setting where the support of the verbs $V$ is different between the two distributions. This means that we can create a generalization problem by restricting a verb $v_{1}$ to never occur in $e_{1}$ but to occur in $e_{2}$ for $p^{train}$, as illustrated by Figure 3.
For $p^{\text{test}}$ we invert the distribution of $e_{1}$, making it a degenerate distribution with single-point support for the verb that is not in $e_{1}$ for $p^{\text{train}}$, while $e_{2}$ has the same distribution as in $p^{\text{train}}$. When generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario, a model will be exposed to all verbs in all permissible syntactic positions in some embedded sequence, but $v_{1}$ will never be seen in $e_{1}$ during training, while $e_{1}$ in $p^{\text{test}}$ always contains $v_{1}$.
@@
The main contribution of this work is to demonstrate that the difficulty of generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario depends on the entropy $\mathrm{H}$ of the distribution of verbs in $e_{2}$ for $p^{\text{train}}$:
@@ 3.1 Increasing Entropy
In this work, we experiment with two different approaches to increasing $H$, as illustrated in Figure 2. In the following sections we formalize these approaches, and in Section 4 we demonstrate how the performance of neural architectures scales for each type of increment.
@@ 3.2 Data Generation
In our experiments, we use the standard conjunctive phrases in SCAN $|C|=2$. Furthermore, we let the two conjunctive phrases take the embedded sentences in the opposite order: if $v_{1}$ always occur in $e_{2}$ for $c_{1}$, then $v_{1}$ will always occur in $e_{1}$ for $c_{2}$.
@@ 4 Experiments
We validate our framework by evaluating the three most common sequence-to-sequence architectures: the Transformer, the RNN, and the CNN. We also run a model with built-in compositional priors to showcase how a model that encodes verb equivariance can solve the task even when $\mathrm{H}$ is low. We have two main experimental settings, corresponding to the two methods for increasing $\mathrm{H}$ outlined in Section 3.1.1 and Section 3.1.2. We also conduct two supporting experiments. For all experimental settings, $\mathrm{H}=0$ constitutes the most difficult generalization problem, and $\mathrm{H}=\log_{2}|V|=3$ constitutes the easiest problem. We average and report results from five differently seeded runs. The following sections provide details about our experimental setup. [^3] All code related to our experiments will be released upon publication at anonymous-url.
@@ 4.1 Models
Lastly, we experiment with a model that has built-in architectural priors. We use the encoder-decoder from Gordon et al. (2019), which enforces the verb equivariance found in SCAN using a cyclic group. This group is used in the encoding of the input sequences in a way that pools the representation of any verb with the representation of the other verbs, even when they are not present in the input. This is a strong structural prior that makes the model able to solve the original SCAN splits efficiently. However, this approach is dataset-dependent, as the elements of the cyclic group must be defined manually. We use this model primarily as a sanity check for our experiments, but also as a comparison for the other architectures, comparing their efficiency to an idealized case.
@@ 4.2 Experiment 1: Vertical scaling
For Experiment 1, $p^{\text{test}}$ has a degenerate distribution for $e_{1}$, where $v_{1}\in V$ always occur, while $e_{2}$ has a uniform distribution over $V$. For $p^{\text{train}}$, $e_{1}$ has a uniform distribution over $V\setminus\{v_{1}\}$, and $e_{2}$ is a mixture of a degenerate distribution at $v_{1}$ and a uniform distribution over $V\setminus\{v_{1}\}$, where we increase the total samples drawn from the uniform while decreasing the number drawn from the degenerate distribution. By doing this we increase $\mathrm{H}$ as described in section Section 3.1.1. We refer to this increase in $\mathrm{H}$ as vertical scaling, as the distribution over the verbs increases towards the uniform. The models are trained on 6 000 unique samples and evaluated on 7 056 samples for all values of $\mathrm{H}$.
@@
Figure 5: Results from the experiment 1 & 2. The accuracy and standard deviations are from five seeds.
@@ 4.4 Effect of Number of Unique Samples
As $H$ describes the relative distribution of verbs in $e_{2}$, it is possible to construct datasets of varying sizes that have the same $\mathrm{H}$. This can be achieved by increasing the total volume of unique samples while maintaining the same relative distribution. To verify that performance scales with $\mathrm{H}$ independently of the number of unique samples seen during training, we create datasets that have the same entropy as in the previous experiments, but where we vary the number of total unique samples. We experiment with three different sample sizes: 3 000, 4 000, and 6 000, using the vertical scaling setting. All models run for the same number of optimization steps for all sample sizes.
@@ 4.5 Effect of Positional Embedding Type
In our experiments, there is an intuitive connection between the generalization from $p^{\text{train}}$ to $p^{\text{test}}$ and the ability to encode verb equivariance. As the verbs always occur on the first token position of $e_{1}$ and $e_{2}$, the positional information is central to both of our experiments.
To assess the effect of positional information, we show how the choice of positional encoding scheme affects performance in the vertical scaling setting. In addition to the original absolute position embeddings from Vaswani (2017), we experiment with embeddings that encode relative positional information through disentangled attention (He et al., 2021) and rotary position embeddings (RoPE; Su et al., 2024). In this experiment, we focus on the Transformer. Details on hyperparameter selection can be found in Section B.4

### v2

@@ 1 Introduction
Human language is characterized by its combinatorial properties in syntax and semantics (Hadley, 1994). For example, when someone knows what it means for Person A to see Person B, they can readily generalize to what it means for Person B to see Person A, without having been exposed to examples where Person B is used in the subject position (Fodor and Pylyshyn, 1988). This type of bidirectional systematic generalization has long been considered a key feature of human language understanding (Chomsky, 1957), and is closely related to the compositionality of natural language semantics, a contentious but widely studied topic in both formal semantics (van Bethem et al., 1991) and NLP (McCurdy et al., 2024).
@@
Early critics of connectionism, such as Fodor and Pylyshyn (1988), argued that systematic generalization was incompatible with continuous representations. More than thirty years since this critique was initially presented, researchers still agree that achieving robust compositional behavior in neural networks remains an open problem, as shown by a recent survey by McCurdy et al. (2024). Research continues to demonstrate that current models struggle with systematic generalization (Lake and Baroni, 2018; Hupkes et al., 2020; Dziri et al., 2024; Wold et al., 2024; Mirzadeh et al., 2024).
Some of the limitations of current models can be alleviated by implementing compositional priors, either in the architecture (Gordon et al., 2019) or through the training procedure (Lake and Baroni, 2023), but it remains unclear whether architectural constraints cause the limitations or whether they arise due to distributional properties of the data used for training.
@@
Figure 2: A schematic overview of the two approaches to increasing the entropy level. Top: Vertical scaling by distribution mixing. Bottom: Horizontal scaling by incrementing the support of the distribution.
In this paper, we ask the following question: Can neural networks display systematic out-of-distribution generalization without any built-in priors? To answer this, we take a data-centric approach. Specifically, we show how the degree of systematic generalization of common architectures relates to the information entropy of the distribution of component parts found in the training data.
Contrary to previous work, we find that popular sequence-to-sequence models used in NLP are capable of systematic generalization, even without built-in architectural priors. However, this is only the case when the training dataset contains high entropy. On a series of sequence-to-sequence datasets based on a modified version of the SCAN (Lake and Baroni, 2018) grammar, we show that model performance scales with the entropy of the training dataset, as illustrated in Figure 1, thereby creating a link between systematic generalization and information efficiency.
@@ 2 Background
Systematic generalization is typically described as a robust mode of composition, where a system is able to combine parts of its training data in a way that enables generalization to novel combinations. As such, systematic generalization relates to compositionality, a topic which has both a long tradition and multiple formal definitions, most prominently in formal semantics (van Bethem et al., 1991; Pagin and Westersthl, 2010). These definitions, however, often prove too restrictive when working with natural languages. As a result, researchers frequently introduce problem-specific relaxations, leading to ambiguity over time regarding the exact meaning of terms such as systematic generalization and compositionality.
Despite there being some ambiguity regarding the formal definitions, recent work shows that there is high agreement among researchers on the following, more informal definition of what constitutes compositional behavior:
@@
“When a model receives an input $I$ that humans conceive as composed of component parts, if the model produces correct outputs for those parts (in isolation or in other combinations), then it will also produce a correct output for $I$ (McCurdy et al., 2024, p. 9324).”
@@ 2.1 Systematic Generalization
Following earlier works, we will refer to the systematic generalization abilities of a model as the systematicity of that model. Our view of systematicity is based on Hadley (1994), who in contrast to earlier work by Fodor and Pylyshyn (1988) distinguished between degrees of systematicity. In the following paragraphs, we present and define these degrees under the assumption that we are considering a system tasked to generalize from a training corpus to a held-out test corpus: [^1] Hadley (1994) also discusses an intermediate degree between weak and strong systematicity, so-called quasi-systematicity, but we leave this out for brevity, as it does not relate directly to our use case.
Definition 2 (Weak systematicity)
@@
Definition 3 (Strong systematicity)
Now, a system displays strong systematicity if it (i) displays weak systematicity, meaning that it can generalize in-distribution, and (ii) it can process novel embedded sentences containing previously learned component parts that occur in syntactic positions where they do not appear in the training data, neither in embedded sentences nor in non-embedded sentences, and (iii) a significant fraction of the possible combinations of component parts and their permissible syntactic position should be left out, based on the observation that humans learn grammars based on largely incomplete data Pinker (1989). In this context, an embedded sentence can be thought of as a syntactic constituent in the overall sentence. For example, if a system can process the following novel sentence Mathilde knows Sigurd and Sigurd knows Olav, which has the sentences Mathilde knows Sigurd and Sigurd knows Olav as embedded sentences, without having seen the name Mathilde previously used in the subject position in a sentence with constituents nor in a sentence with no constituents, then it is strongly systematic w.r.t to how it can generalize to the novel use of the name Mathilde. On the other hand, if the system can only process this novel sentence if it has seen Mathilde being used in the same syntactic position, for example in the sentence Mathilde knows Olav, then it is not strongly systematic.
@@
In this work, we focus on a relaxed version of strong systematicity as defined in Definition 3. We relax requirement (ii) from the original definition so that we only work with samples that have embedded sentences. As we show in the following section, this makes it easier to quantify the level of systematicity required for a given generalization problem, which is the main focus of this work. Furthermore, we also relax (ii) so that it allows component parts of the test set to occur at the same syntactic position in some embedded sentence in the train set, but not in an embedded sentence that occurs at the same position in the overall sequence.
In order to compare models with and without compositional priors from the literature, we base our experiments on the SCAN grammar (Lake and Baroni, 2018). This grammar has few permissible syntactic positions for each part of speech, making the original requirement from Hadley (1994) too restrictive with respect to the overall sample size needed by neural networks. In the following section, we formalize our framework and elaborate on these choices.
@@ 3 Measuring Systematicity
In this work, we demonstrate that sequence-to-sequence models can generalize to embedded sentences where component parts occur at novel syntactic positions in the overall sequence. For our experiments, we construct datasets where each dataset sample contains an input sequence $x\in X$ and an output sequence $y\in Y$, based on a modified version of the SCAN grammar. We discuss the specifics of this dataset generation procedure, including details about the context-free grammar that generates $x$, in Section 3.5. In this section, we define a framework for measuring the degree of systematicity displayed by a systematic generalization problem.
3.1 Task Description
Throughout this paper, $x$ comprises two sentences conjoined by a conjunctive phrase. [^2] In contrast to the original SCAN, we do not have sentences without a conjunctive phrase. Consequently, $x$ can be modeled as a triple, with a conjunction $\operatorname{c}\in C$, and the two embedded sentences $e_{1},e_{2}\in\Sigma^{*}$ as arguments, $x=(e_{1},\operatorname{c},e_{2})$. The component parts of both embedded sentences are drawn from the same vocabulary $\Sigma=V\cup A$, where $V$ is a set of verbs and $A$ is a set of adverbial phrases, e.g.: jump, run $\in V$, twice $\in A$, and and $\in C$. There exists a deterministic procedure $\varphi$ for transforming $x$ into $y$:
@@ 3.1 Task Description
The goal of the described task is to learn a function $f$ that approximates $\varphi$, which requires modeling the conditional distribution of output sequences given input sequences in the training data, i.e. $p^{\text{train}}(y\mid x)$. See Table 1 for more examples generated by our modified SCAN grammar.
3.2 Component Distributions
Since we are interested in systematic generalization, we also need to define the distribution of component parts in $e_{1}$ and $e_{2}$, and we are specifically interested in the case where the distribution of $e_{1}$ and $e_{2}$ is different in the training and test data, but they share a similar distribution when marginalized over position, which in this case refers to which side of the conjunctive phrase they appear in.
In SCAN (Appendix A), $e_{1}$ and $e_{2}$ always contain one verb in $V$ and zero or more adverbial phrases from $A$ that control how many times that verb should be repeated, and in what direction it is executed. To create a systematic generalization problem that adheres to the definition given in Section 2.1, it is necessary to set the support of $e_{1}$ and $e_{2}$ to be different in the training and test distributions. For simplicity, we limit our study to focus on the setting where the support of the verbs $V$ is different between the two distributions. This means that we can create a generalization problem by restricting a verb $v_{1}$ to never occur in $e_{1}$ but to occur in $e_{2}$ for $p^{train}$, as illustrated by Figure 3.
For $p^{\text{test}}$ we invert the distribution of $e_{1}$, making it a degenerate distribution with single-point support for the verb that is not in $e_{1}$ for $p^{\text{train}}$, while $e_{2}$ has the same distribution as in $p^{\text{train}}$. When generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in this scenario, a model will be exposed to all verbs in all permissible syntactic positions in some embedded sentence, but $v_{1}$ will never be seen in $e_{1}$ during training, while $e_{1}$ in $p^{\text{test}}$ always contains $v_{1}$.
@@ 3.2 Component Distributions
[…]
Figure 4: Probabilities in the degenerate case for $e_{2}$
3.3 Systematicity as Entropy
The main contribution of this work is to demonstrate that the difficulty of generalizing from $p^{\text{train}}$ to $p^{\text{test}}$ in the scenario described above depends on the entropy $\mathrm{H}$ of the distribution of verbs in $e_{2}$ for $p^{\text{train}}$:
@@ 3.4 Increasing Entropy
In this work, we experiment with two different approaches to increasing $H$, as illustrated in Figure 2. In the following sections, we formalize these approaches, and in Section 4 we demonstrate how the performance of neural architectures scales for each type of increment.
@@ 3.5 Data Generation
In our experiments, we use the standard conjunctive phrases in SCAN $|C|=2$. Furthermore, we let the two conjunctive phrases take the embedded sentences in the opposite order: if $v_{1}$ always occurs in $e_{2}$ for $c_{1}$, then $v_{1}$ will always occur in $e_{1}$ for $c_{2}$.
@@ 4 Experiments
We validate our framework by evaluating the three most common sequence-to-sequence architectures: the Transformer, the RNN, and the CNN. We also run a model with built-in compositional priors to showcase how a model that encodes verb equivariance can solve the task even when $\mathrm{H}$ is low. We have two main experimental settings, corresponding to the two methods for increasing $\mathrm{H}$ described in Section 3.4.1 and Section 3.4.2. We also conduct a supporting experiment to control for a potential confounding effect resulting from an increased sample size, and an experiment that estimates the influence of different position encoding types. For all experimental settings, $\mathrm{H}=0$ constitutes the most difficult generalization problem, and $\mathrm{H}=\log_{2}|V|=3$ constitutes the easiest problem. We average and report results from five differently seeded runs. The following sections provide details about our experimental setup. [^3] Data generation scripts and code to reproduce our experiments can be found at https://github.com/ltgoslo/systematicity-entropy.
@@ 4.1 Models
Lastly, we experiment with a model that has built-in architectural priors. We use the encoder-decoder from Gordon et al. (2019), which enforces the verb equivariance found in SCAN using a cyclic group. This group is used in the encoding of the input sequences in a way that pools the representation of any verb with the representation of the other verbs, even when they are not present in the input. This is a strong structural prior that enables the model to solve the original SCAN splits efficiently. However, this approach is dataset-dependent, as the elements of the cyclic group must be defined manually. We use this model primarily as a sanity check for our experiments, but also as a comparison for the other architectures, comparing their efficiency to an idealized case.
@@ 4.2 Experiment 1: Vertical scaling
For Experiment 1, $p^{\text{test}}$ has a degenerate distribution for $e_{1}$, where $v_{1}\in V$ always occur, while $e_{2}$ has a uniform distribution over $V$. For $p^{\text{train}}$, $e_{1}$ has a uniform distribution over $V\setminus\{v_{1}\}$, and $e_{2}$ is a mixture of a degenerate distribution at $v_{1}$ and a uniform distribution over $V\setminus\{v_{1}\}$, where we increase the total samples drawn from the uniform while decreasing the number drawn from the degenerate distribution. By doing this, we increase $\mathrm{H}$ as described in section Section 3.4.1. We refer to this increase in $\mathrm{H}$ as vertical scaling, as the distribution over the verbs increases towards the uniform. The models are trained on 6 000 unique samples and evaluated on 7 056 samples for all values of $\mathrm{H}$.
@@
Figure 5: Results from the Experiments 1 & 2. The accuracy and standard deviations are from five seeds.
@@ 4.4 Effect of Number of Unique Samples
As $\mathrm{H}$ describes the relative distribution of verbs in $e_{2}$, it is possible to construct datasets of varying sizes that have the same $\mathrm{H}$. This can be achieved by increasing the total volume of unique samples while maintaining the same relative distribution. To verify that performance scales with $\mathrm{H}$ independently of the number of unique samples seen during training, we create datasets that have the same entropy as in the previous experiments, but where we vary the number of total unique samples. We experiment with three different sample sizes: 3 000, 4 000, and 6 000, using the vertical scaling setting. All models run for the same number of optimization steps for all sample sizes.
@@ 4.5 Effect of Positional Embedding Type
In our experiments, there is an intuitive connection between the generalization from $p^{\text{train}}$ to $p^{\text{test}}$ and the ability to encode verb equivariance. This is exemplified by Gordon et al. (2019), which achieves high performance on the original SCAN tasks precisely by enforcing verb equivariance in the model architecture explicitly.
As the verbs always occur on the first token position of $e_{1}$ and $e_{2}$, positional information is central to our experiments. To estimate the effect of positional information, we show how the choice of positional encoding scheme affects performance in the vertical scaling setting. In addition to the original absolute position embeddings from Vaswani (2017), we experiment with embeddings that encode relative positional information through disentangled attention (He et al., 2021) and rotary position embeddings (RoPE; Su et al., 2024). In this experiment, we focus on the Transformer. Details on hyperparameter selection can be found in Section B.4